COMBINED ANALYSIS OF THE Ã^{2}E′′ −X̃^{2}E′ GAS-PHASE EXCITATION AND EMISSION SPECTRA OF Ag_{3}

\underline{ILIAS SIOUTIS}, VADIM STAKHURSKY, RUSSELL M. PITZER, and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Metal clusters are of considerable interest as they bridge the atomic and solid states. Among the heavier homonuclear coinage metals is silver trimer Ag_{3}, a particularly interesting molecule in that although the orbitally degenerate X̃^{2}E′ and Ã^{2}E′′ D_{3h} states can undergo Jahn-Teller (JT) distortion to C_{2v} states, spin-orbit splitting may partially quench this distortion. Both laser-induced fluorescence (LIF) and dispersed fluorescence (DF) spectra have been reported for Ag_{3}. Preliminary JT analyses for the Ã^{2}E′′ and X̃^{2}E′ states have been based, respectively, on the LIF and DF data. However the interpretation has been subject to some ambiguity. High level ab-initio calculations have been performed with the express purpose of estimating the Jahn-Teller constants and spin-orbit effects in both states. These theoretically predicted molecular parameters initiated the Jahn-Teller simulations of the LIF and DF spectra, leading to an overall understanding of the spin-vibronic structure in each case.